OpenStudy (anonymous):

prove euler's identities: sin z =((e^iz)-(e^-iz))/2i and cos z = ((e^iz)+(e^-iz))/2

OpenStudy (anonymous):

e^(iz)=cos(z)+isin(z) ...........1 e^(-iz)=cos(z)-isin(z) ..........2 subtract (2) from (1) you will get it.

OpenStudy (anonymous):

by subtraction you will get sin z =((e^iz)-(e^-iz))/2i and by addition of (1) and (2) you will have cos z = ((e^iz)+(e^-iz))/2

OpenStudy (anonymous):

thanks